​Find the standard deviation of the following data. Round your answer to one decimal place.


Find the standard deviation of the following data. Round your answer to one decimal place. 

x012345
P(X=x)0.20.10.10.20.20.2
Answer #1

Solution:

X P(X) X*P(X) X^2*P(X)
0 0.2 0 0
1 0.1 0.1 0.1
2 0.1 0.2 0.4
3 0.2 0.6 1.8
4 0.2 0.8 3.2
5 0.2 1.7 5.5
Sum = 1 3.4 11

Mean :

\mu = \sum X * P(X)

= 3.4

Standard deviation:

\sigma = \sqrt{} \sum X2 * P(X) - (\mu)2

= \sqrt{} 11 - ( 3.4 )2

= \sqrt{} -0.56

= 0.7